OpenStudy (anonymous):

what is 2AND 4 /20=2 AND BLANK 40

OpenStudy (xishem):

\[2\frac{4}{20}=2\frac{x}{40}\]The 2s will cancel each other out...\[\frac{4}{20}=\frac{x}{40}\]Multiply both sides by 40 to isolate x...\[40(\frac{4}{20})=x \rightarrow x=8\]\[2\frac{4}{20}=2\frac{8}{40}\]

OpenStudy (anonymous):

THANKS /XISHEM